# Optimal Frames for Phase Retrieval from Edge Vectors of Optimal Polygons

### v2

@@ 3.1. The relationship between tight frames and convex polygons in $\mathbb{R}^{2}$
Assume that $\mathcal{A}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset\mathcal{T}$, where $\mathcal{T}$ is defined in (1.9). Let $E=\{\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}\}$, where each $\boldsymbol{e}_{j}=f(\boldsymbol{a}_{j})$, where $f$ is defined in (1.10). Then we have the following results:
@@
    The set $\mathcal{A}$ is irreducible if and only if the vectors in $E$ are nonzero and have pairwise distinct directions.
@@
    The set $\mathcal{A}$ is a tight frame if and only if $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$.
  3. (iii)
    The set $\mathcal{A}$ is an irreducible tight frame if and only if $E$ is the edge set of a strictly convex $m$-gon ${P}\in\mathcal{P}_{m}$.
  4. (iv)
    Assume that the set $\mathcal{A}$ forms a tight frame. Let $\boldsymbol{A}=[\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}]^{T}$. If $\mathcal{A}$ is irreducible, then
@@
    where $P$ is the strictly convex $m$-gon defined in (iii). If $\mathcal{A}$ is reducible, then the condition number of $\boldsymbol{A}$ is not the minimum value attainable among all $m\times 2$ matrices.
@@
(i) We prove the equivalence in two directions. Assume that the set $\mathcal{A}$ is irreducible. First, we show that vectors in $E$ are nonzero. It is important to note that $\boldsymbol{e}_{j}=f(\boldsymbol{a}_{j})$ is the zero vector if and only if $\boldsymbol{a}_{j}$ is the zero vector. Since $\mathcal{A}$ is irreducible, by definition, it does not contain the zero vector. Consequently, given this property of $f$, the set $E$ also does not contain the zero vector. Next, we show that vectors in $E$ have pairwise distinct directions. Suppose, for contradiction, that there exist distinct vectors $\boldsymbol{e}_{i}=f(\boldsymbol{a}_{i})$ and $\boldsymbol{e}_{j}=f(\boldsymbol{a}_{j})$ (with $i\neq j$) that have the same direction. This implies that their associated directional parameters are equal, i.e., $\phi_{j}=\phi_{i}$. This condition is, in turn, equivalent to $\boldsymbol{a}_{j}=\lambda\boldsymbol{a}_{i}$ for some nonzero scalar $\lambda\in\mathbb{R}$. However, this directly contradicts the definition of an irreducible set $\mathcal{A}$, which requires that no two distinct vectors are scalar multiples of each other. Hence, we conclude that if $\mathcal{A}$ is irreducible, then the vectors in $E$ are nonzero and have pairwise distinct directions.
Conversely, if the vectors in $E$ are nonzero and have pairwise distinct directions, the irreducibility of $\mathcal{A}$ can be established by an analogous argument.
(ii) We define
$a=\sum_{j=1}^{m}t_{j}^{2}\cos^{2}{\phi_{j}},\quad b=\sum_{j=1}^{m}t_{j}^{2}\sin^{2}{\phi_{j}},\quad\text{and}\quad c=\sum_{j=1}^{m}t_{j}^{2}\cos{\phi_{j}}\sin{\phi_{j}}.$
A direct computation shows that $\sum_{j=1}^{m}\boldsymbol{e}_{j}=(a-b,2c)^{T}$ and
$\sum_{j=1}^{m}\boldsymbol{a}_{j}\boldsymbol{a}_{j}^{T}=\begin{pmatrix}\sum_{j=1}^{m}t_{j}^{2}\cos^{2}{\phi_{j}}&\sum_{j=1}^{m}t_{j}^{2}\cos{\phi_{j}}\sin{\phi_{j}}\\ \sum_{j=1}^{m}t_{j}^{2}\cos{\phi_{j}}\sin{\phi_{j}}&\sum_{j=1}^{m}t_{j}^{2}\sin^{2}{\phi_{j}}\\ \end{pmatrix}=\begin{pmatrix}a&c\\ c&b\\ \end{pmatrix}$. (3.1)
Therefore, $\{\boldsymbol{a}_{1},\dots,\boldsymbol{a}_{m}\}$ is a tight frame if and only if $a=b$ and $c=0$, which is equivalent to $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$.
(iii) Assume that $\{\boldsymbol{a}_{1},\dots,\boldsymbol{a}_{m}\}$ is an irreducible tight frame. It follows from (i) and (ii) that $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$ and $\boldsymbol{e}_{1},\dots,\boldsymbol{e}_{m}$ have pairwise distinct directions. Since $\boldsymbol{e}_{j}\neq\boldsymbol{0}$ for all $j\in[m]$, according to Corollary 2.1, there exists a strictly convex $m$-gon P{P} with the edge set $E$.
Assume that $E$ is the edge set of a strictly convex $m$-gon ${P}\in\mathcal{P}_{m}$. By Corollary 2.1 we have $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$, and $\boldsymbol{e}_{1},\dots,\boldsymbol{e}_{m}$ are nonzero vectors with pairwise distinct directions. If follows from (i) and (ii) that $\{\boldsymbol{a}_{1},\dots,\boldsymbol{a}_{m}\}$ is an irreducible tight frame.
(iv) Since $\mathcal{A}$ is a tight frame, by Lemma 2.1 we have
@@
If $\mathcal{A}$ is irreducible, it follows from (iii) that $E$ is the edge set of a strictly convex $m$-gon ${P}\in\mathcal{P}_{m}$. Combining (3.4) with Lemma 2.4, we obtain $\beta_{\boldsymbol{A}}=\frac{1}{\sqrt{1-2\cdot r({P})}}$.
We next prove that the condition number of $\boldsymbol{A}$ is not minimal if $\mathcal{A}$ is reducible. Assume $\mathcal{A}$ is reducible. Then the set $E$ contains zero vectors or vectors that have the same direction. By deleting zero vectors from $E$ and replacing all vectors in the same direction with their sum, we obtain a new set $\widehat{E}=\{\widehat{\boldsymbol{e}}_{1},\dotsc,\widehat{\boldsymbol{e}}_{k}\}$ with $k<m$. The vectors in $\widehat{E}$ are nonzero and have pairwise distinct directions. Note that $\sum_{j=1}^{k}\widehat{\boldsymbol{e}}_{j}=\sum_{j=1}^{m}\boldsymbol{e}_{j}=0$. By Corollary 2.1, there exists a strictly convex $k$-gon $\widehat{P}\in\mathcal{P}_{k}$ with the edge set $\widehat{E}$. Note that $\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}=\sum_{j=1}^{k}\|\widehat{\boldsymbol{e}}_{j}\|_{2}$ and $\sum_{j=1}^{m}|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|=\sum_{j=1}^{k}|\langle\widehat{\boldsymbol{e}}_{j},\boldsymbol{u}\rangle|$ for any $\boldsymbol{u}\in\mathbb{S}^{1}$. Combining with (3.4) and Lemma 2.4, we have
@@
Let $Q\in\mathcal{P}_{m}$ be an optimal strictly convex $m$-gon. Denote the edge set of $Q$ by $\{\boldsymbol{q}_{1},\ldots,\boldsymbol{q}_{m}\}$. Let $\boldsymbol{B}=[f^{-1}(\boldsymbol{q}_{1}),\ldots,f^{-1}(\boldsymbol{q}_{m})]^{T}\in\mathbb{R}^{m\times 2}$, where $f(\cdot)$ is defined in (1.10). According to (iii), the set $\{f^{-1}(\boldsymbol{q}_{1}),\ldots,f^{-1}(\boldsymbol{q}_{m})\}$ forms an irreducible tight frame. Then, we have
@@
Recall that any optimal convex m-gon is strictly convex [25]. Since $\widehat{P}\in\mathcal{P}_{k}$ is a convex $m$-gon but not strictly convex (it has a flat angle), it cannot be optimal. Hence, we have $r(\widehat{P})>r({Q})$.
@@
Combining with (3.5) and (3.6), we have $\beta_{\boldsymbol{A}}>\beta_{\boldsymbol{B}}$. Therefore, the condition number of $\boldsymbol{A}$ is not minimal. This completes the proof.
@@
(i) Assume that $\mathcal{A}$ is an optimal tight frame, i.e., the matrix $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}]^{T}$ has the minimal condition number and its rows form a tight frame. It follows from Lemma 3.1 (iv) that $\mathcal{A}$ is irreducible. Then Lemma 3.1 (iii) shows that the set $\{f(\boldsymbol{a}_{1}),\dotsc,f(\boldsymbol{a}_{m})\}$ is the edge set of a strictly convex $m$-gon ${P}\subset\mathbb{R}^{2}$, and we have $\beta_{{\boldsymbol{A}}}=\frac{1}{\sqrt{1-2\cdot r(P)}}$. We prove that $P$ is optimal by contradiction. Suppose, to the contrary, that $P$ is not optimal. Then there exists a strictly convex $m$-gon ${Q}\subset\mathbb{R}^{2}$ such that $r(Q)<r(P)$. Denote the edge vectors set of $Q$ by $\{\boldsymbol{q}_{1},\ldots,\boldsymbol{q}_{m}\}$. Let $\boldsymbol{B}=[f^{-1}(\boldsymbol{q}_{1}),\ldots,f^{-1}(\boldsymbol{q}_{m})]^{T}\in\mathbb{R}^{m\times 2}$. By Lemma 3.1 (iii) we have $\beta_{\boldsymbol{B}}=\frac{1}{\sqrt{1-2\cdot r({Q})}}$. Since $r(Q)<r(P)$, we have $\beta_{\boldsymbol{B}}<\beta_{\boldsymbol{A}}$. This contradicts with the fact that $\boldsymbol{A}$ has the minimal condition number. Hence, $P$ is optimal.
(ii) Assume that $\{f(\boldsymbol{a}_{1}),\dotsc,f(\boldsymbol{a}_{m})\}$ is the edge set of an optimal $m$-gon $P\in\mathcal{P}_{m}$. Then $P$ is strictly convex. According to Lemma 3.1 (iii) and (iv), the set $\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}$ is an irreducible tight frame, and we have $\beta_{{\boldsymbol{A}}}=\frac{1}{\sqrt{1-2\cdot r(P)}}$. We prove that $\boldsymbol{A}$ has the minimal condition number by contradiction. Assume that $\beta_{\boldsymbol{B}}<\beta_{\boldsymbol{A}}$, where $\boldsymbol{B}\in\mathbb{R}^{m\times 2}$ is a matrix with the minimal condition number and its rows form a tight frame. By Lemma 3.1 (iii), the rows of $\boldsymbol{B}$ is irreducible, and there is a strictly convex $m$-gon ${Q}\subset\mathbb{R}^{2}$ such that $\beta_{{\boldsymbol{B}}}=\frac{1}{\sqrt{1-2\cdot r(Q)}}$. Since $\beta_{\boldsymbol{B}}<\beta_{\boldsymbol{A}}$, we have $r(Q)<r(P)$. This contradicts with $P$ being an optimal convex $m$-gon. Hence, $\boldsymbol{A}$ has the minimal condition number, and its rows form a tight frame.
3.3. Proof of Corollary 1.1
Proof of Corollary 1.1.
By Lemma 2.2 we have
4. Proofs of Theorem 1.3 and Theorem 1.4
4.1. Characterization of optimal convex polygons in $\mathbb{R}^{2}$ for the perimeter-maximizing isodiametric problem
We begin by recalling the definition of a Reuleaux polygon, which will be used in the sequel.
@@
Let $P$ be an optimal convex $m$-gon, where $m\geq 3$ is an integer with an odd factor. By Theorem 4.2 and Theorem 4.1 (i), $P$ is inscribed in a Reuleaux polygon $R$ with $r$ boundary arcs, where $r$ is an odd integer with $3\leq r\leq m$. Label the vertices of $R$ as $\boldsymbol{w}_{1},\dotsc,\boldsymbol{w}_{r}$ in counterclockwise order, and extend the indexing cyclically by setting $\boldsymbol{w}_{r+i}=\boldsymbol{w}_{i}$ for each $i\in[r]$. Define $h=\frac{r-1}{2}$. Then for each $i\in[r]$, the vertex $\boldsymbol{w}_{i}$ is the the center of the arc ¿ $\boldsymbol{w}_{{i+h}}\boldsymbol{w}_{{i+h+1}}$ on the boundary of $R$. This can be seen as follows. By Theorem 4.1 (ii), we can start at $\boldsymbol{s}_{1}:=\boldsymbol{w}_{1}$ and construct a star polygon $S$ by sequentially connecting pairs of vertices in $R$ at maximal distance in the counterclockwise direction. Let $\boldsymbol{s}_{i}$ denote the $i$-th vertex of $R$ visited during this construction. According to [25, Page 256], the vertices $\boldsymbol{s}_{1},\boldsymbol{s}_{3},\dotsc,\boldsymbol{s}_{r},\boldsymbol{s}_{2},\boldsymbol{s}_{4},\dotsc,\boldsymbol{s}_{r-1}$ are ordered counterclockwise. Comparing this with the labeling $\boldsymbol{w}_{1},\ldots,\boldsymbol{w}_{r}$, we have
@@
Figure 3. One of the optimal dodecagon as in Figure 2 (a2), with vertices $\boldsymbol{v}_{1},\dotsc,\boldsymbol{v}_{12}$. The corresponding Reuleaux polygon consists of $9$ circular arcs with vertices $\boldsymbol{v}_{c_{1}},\dotsc,\boldsymbol{v}_{c_{9}}$ indicated in red dots. The dashed lines connect pairs of vertices at maximal distance.
@@
Assume that $P\in{\mathcal{E}}_{m}$. It follows from Theorem 4.2 that $P$ is equilateral. We now prove that $P$ satisfies (4.1). Assume without loss of generality that $\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}$ are arranged counterclockwise. Let $\boldsymbol{v}_{1},\dotsc,\boldsymbol{v}_{m}$ be the vertices of $P$ such that $\boldsymbol{e}_{j}=\boldsymbol{v}_{j+1}-\boldsymbol{v}_{j}$ for each $j\in[m-1]$ and $\boldsymbol{e}_{m}=\boldsymbol{v}_{1}-\boldsymbol{v}_{m}$. For each $i\in[m]$, we denote $\boldsymbol{v}_{m+i}=\boldsymbol{v}_{i}$, and let $\theta_{i}$ be the counterclockwise angle from $\boldsymbol{e}_{i-1}$ to $\boldsymbol{e}_{i}$, where $\boldsymbol{e}_{0}:=\boldsymbol{e}_{m}$. Then we have $\alpha_{1}=0$ and $\alpha_{i}=\alpha_{i-1}+\theta_{i}$ for each $i\in\{2,\ldots,m\}$. Moreover, since $P$ is strictly convex, we have $\theta_{i}\in(0,\pi)$, $\forall i\in[m]$.
@@
4.2. Proof of Theorem 1.4
The following lemma is useful in the proof of Theorem 1.4, which is essentially proved in [32] (see equation (2.2) and (3.16) in [32]).
Lemma 4.2.
$(L_{\boldsymbol{A}})^{2}\leq\frac{1}{2}\sum_{j=1}^{m}\|\boldsymbol{a}_{j}\|_{2}^{2}-\frac{1}{2m\cdot\sin\frac{\pi}{2m}}\sum_{j=1}^{m}\|\boldsymbol{a}_{j}\|_{2}^{2},$
where $L_{\boldsymbol{A}}$ is the optimal lower Lipschitz constant of the map $\Phi_{\boldsymbol{A}}$ defined in (1.1).
Now we can present a proof of Theorem 1.4.
Proof of Theorem 1.4.
Combining with Lemma 4.2, we derive that

### v3

@@ 2.3. The perimeter-maximizing isodiametric problem
Assume that $\mathcal{A}=\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}\subset\mathcal{T}$, where $\mathcal{T}$ is defined in (1.7). Let $E=\{\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}\}$, where each $\boldsymbol{e}_{j}=f(\boldsymbol{a}_{j})$ with $f$ defined in (3.1). Then the following hold:
@@
    $\mathcal{A}$ is a tight frame consisting of nonzero vectors with pairwise distinct directions if and only if $E$ is the edge set of a strictly convex $m$-gon ${P}\in\mathcal{P}_{m}$.
@@
    Assume that $\mathcal{A}$ forms a tight frame. If the vectors in $\mathcal{A}$ are nonzero and have pairwise distinct directions, then
@@
    where $\boldsymbol{A}=[\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}]^{T}$ and $P$ is the strictly convex $m$-gon defined in (i); otherwise, the condition number of $\boldsymbol{A}$ is not the minimum value attainable among all $m\times 2$ matrices.
@@
(i) Recall that $\{\boldsymbol{a}_{1},\dots,\boldsymbol{a}_{m}\}$ is a tight frame if and only if $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$ [19, Lemma 4.1]. Note that $\boldsymbol{e}_{j}=f(\boldsymbol{a}_{j})=\boldsymbol{0}$ if and only if $\boldsymbol{a}_{j}=\boldsymbol{0}$. Also note that for any $i\neq j$, the vectors $\boldsymbol{e}_{i}=f(\boldsymbol{a}_{i})$ and $\boldsymbol{e}_{j}=f(\boldsymbol{a}_{j})$ have the same direction if and only if $\boldsymbol{a}_{i}$ and $\boldsymbol{a}_{j}$ have the same direction. Therefore, $\{\boldsymbol{a}_{1},\dots,\boldsymbol{a}_{m}\}$ is a tight frame of nonzero vectors with pairwise distinct directions if and only if $\sum_{j=1}^{m}\boldsymbol{e}_{j}=\boldsymbol{0}$ and $\boldsymbol{e}_{1},\dots,\boldsymbol{e}_{m}$ are nonzero vectors with pairwise distinct directions. By Corollary 2.1, this is further equivalent to $E$ being the edge set of a strictly convex $m$-gon ${P}\in\mathcal{P}_{m}$.
(ii) Since $\mathcal{A}$ is a tight frame, by Lemma 2.2 we have
@@
If the vectors in $\mathcal{A}$ are nonzero and have pairwise distinct directions, it follows from (i) that $E$ is the edge set of a strictly convex $m$-gon ${P}\in\mathcal{P}_{m}$. Combining (3.5) with Lemma 2.5, we obtain $\beta_{\boldsymbol{A}}=\frac{1}{\sqrt{1-2\cdot r({P})}}$.
We next prove that the condition number of $\boldsymbol{A}$ is not minimal whenever $\mathcal{A}$ contains zero vectors or vectors that have the same direction. Assume $\mathcal{A}$ contains such vectors. Then so does the set $E$. By deleting zero vectors from $E$ and replacing all vectors in the same direction with their sum, we obtain a new set $\widehat{E}=\{\widehat{\boldsymbol{e}}_{1},\dotsc,\widehat{\boldsymbol{e}}_{k}\}$ with $k<m$. The vectors in $\widehat{E}$ are nonzero and have pairwise distinct directions. Note that $\sum_{j=1}^{k}\widehat{\boldsymbol{e}}_{j}=\sum_{j=1}^{m}\boldsymbol{e}_{j}=0$. By Corollary 2.1, there exists a strictly convex $k$-gon $\widehat{P}\in\mathcal{P}_{k}$ with the edge set $\widehat{E}$. Note that $\sum_{j=1}^{m}\|\boldsymbol{e}_{j}\|_{2}=\sum_{j=1}^{k}\|\widehat{\boldsymbol{e}}_{j}\|_{2}$ and $\sum_{j=1}^{m}|\langle\boldsymbol{e}_{j},\boldsymbol{u}\rangle|=\sum_{j=1}^{k}|\langle\widehat{\boldsymbol{e}}_{j},\boldsymbol{u}\rangle|$ for any $\boldsymbol{u}\in\mathbb{S}^{1}$. Combining with (3.5) and Lemma 2.5, we have
@@
Recall that any optimal convex m-gon is strictly convex [26]. Since $\widehat{P}\in\mathcal{P}_{k}$ is a convex $m$-gon but not strictly convex (it has a flat interior angle), it cannot be optimal. Hence, there exists a strictly convex $m$-gon $Q\in\mathcal{P}_{m}$ such that $r(\widehat{P})>r({Q})$. Denote the edge set of $Q$ by $\{\boldsymbol{q}_{1},\ldots,\boldsymbol{q}_{m}\}$. Let $\boldsymbol{B}=[f^{-1}(\boldsymbol{q}_{1}),\ldots,f^{-1}(\boldsymbol{q}_{m})]^{T}\in\mathbb{R}^{m\times 2}$, where $f(\cdot)$ is defined in (3.1). According to (i), the set $\{f^{-1}(\boldsymbol{q}_{1}),\ldots,f^{-1}(\boldsymbol{q}_{m})\}$ consists of nonzero vectors with pairwise distinct directions. Thus, we have
@@
Recall that $r(\widehat{P})>r({Q})$. Combining with (3.6) and (3.7), we have $\beta_{\boldsymbol{A}}>\beta_{\boldsymbol{B}}$. Therefore, the condition number of $\boldsymbol{A}$ is not minimal. This completes the proof. ∎
We are now prepared to present the proof of Theorem 3.1.
Proof of Theorem 3.1.
(i) Assume that $\mathcal{A}$ is an optimal tight frame, i.e., the matrix $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{m}]^{T}$ has the minimal condition number and its rows form a tight frame. It follows from Lemma 3.1 (ii) that $\mathcal{A}$ consists of nonzero vectors with pairwise distinct directions. Then Lemma 3.1 (i) shows that the set $\{f(\boldsymbol{a}_{1}),\dotsc,f(\boldsymbol{a}_{m})\}$ is the edge set of a strictly convex $m$-gon ${P}\in\mathcal{P}_{m}$, and Lemma 3.1 (ii) further yields that $\beta_{{\boldsymbol{A}}}=\frac{1}{\sqrt{1-2\cdot r(P)}}$. We prove that $P$ is optimal by contradiction. Suppose, to the contrary, that $P$ is not optimal. Then there exists a strictly convex $m$-gon ${Q}\in\mathcal{P}_{m}$ such that $r(Q)<r(P)$. Denote the edge vector set of $Q$ by $\{\boldsymbol{q}_{1},\ldots,\boldsymbol{q}_{m}\}$. Let $\boldsymbol{B}=[f^{-1}(\boldsymbol{q}_{1}),\ldots,f^{-1}(\boldsymbol{q}_{m})]^{T}\in\mathbb{R}^{m\times 2}$. By Lemma 3.1 (i) and (ii), we have $\beta_{\boldsymbol{B}}=\frac{1}{\sqrt{1-2\cdot r({Q})}}$. Since $r(Q)<r(P)$, we have $\beta_{\boldsymbol{B}}<\beta_{\boldsymbol{A}}$. This contradicts with the fact that $\boldsymbol{A}$ has the minimal condition number. Hence, $P$ is optimal.
(ii) Assume that $\{f(\boldsymbol{a}_{1}),\dotsc,f(\boldsymbol{a}_{m})\}$ is the edge set of an optimal $m$-gon $P\in\mathcal{P}_{m}$. Then $P$ is strictly convex. According to Lemma 3.1 (i) and (ii), the set $\{\boldsymbol{a}_{1},\dotsc,\boldsymbol{a}_{m}\}$ is a tight frame, and we have $\beta_{{\boldsymbol{A}}}=\frac{1}{\sqrt{1-2\cdot r(P)}}$. We prove that $\boldsymbol{A}$ has the minimal condition number by contradiction. Assume that $\beta_{\boldsymbol{B}}<\beta_{\boldsymbol{A}}$, where $\boldsymbol{B}\in\mathbb{R}^{m\times 2}$ is a matrix with the minimal condition number and its rows form a tight frame. By Lemma 3.1 (ii), the row vectors of $\boldsymbol{B}$ are nonzero and have pairwise distinct directions, and there is a strictly convex $m$-gon ${Q}\in\mathcal{P}_{m}$ such that $\beta_{{\boldsymbol{B}}}=\frac{1}{\sqrt{1-2\cdot r(Q)}}$. Since $\beta_{\boldsymbol{B}}<\beta_{\boldsymbol{A}}$, we have $r(Q)<r(P)$. This contradicts with $P$ being an optimal convex $m$-gon. Hence, $\boldsymbol{A}$ has the minimal condition number, and its rows form a tight frame. ∎
4. Characterization of optimal frames and optimal polygons: Proof of Theorem 1.2
The aim of this section is to prove Theorem 1.2, which characterizes all optimal frames in $\mathbb{R}^{2}$ with $m$ vectors when $m\geq 3$ has an odd factor. Our approach continues to draw on tools from the perimeter-maximizing isodiametric problem: we first analyze the structure of all optimal $m$-gons, and the proof of Theorem 1.2 then follows by combining this analysis with Theorem 3.1.
4.1. Characterization of optimal polygons
Let ${\mathcal{E}}_{m}$ denote the collection of all optimal convex $m$-gons in $\mathbb{R}^{2}$. The characterization of optimal polygons $P\in{\mathcal{E}}_{m}$ begins with the introduction of the following problem concerning the discrepancy of roots of unity. This problem has already appeared in (1.8), and here we state it formally.
Problem 4.1.
In the following, we demonstrate a close connection between the above discrepancy problem and the optimal polygons. The proof of Theorem 4.1 is postponed to the end of this section.
Theorem 4.1.
Assume that $m\geq 3$ be an integer with an odd factor. Set $\boldsymbol{\mu}_{j}:=(\cos\frac{j\pi}{m},\sin\frac{j\pi}{m})^{T}$ for each $j\in[m]$. Then we have
  1. (i)
  2. (ii)
Remark 4.1.
For a vector $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dotsc,\varepsilon_{m})^{T}\in\{\pm 1\}^{m}$, we define the shifting operator $\mathcal{S}$ and the reflection operator $\mathcal{R}$ as
  1. (i)
    if $P_{2}$ is obtained from $P_{1}$ by a proper translation and scaling, then $\boldsymbol{\varepsilon}_{2}=\boldsymbol{\varepsilon}_{1}$;
  2. (ii)
  3. (iii)
    if $P_{2}$ is obtained from $P_{1}$ by a reflection with respect to the $y$-axis, then $\boldsymbol{\varepsilon}_{2}=\mathcal{S}(\mathcal{R}(\boldsymbol{\varepsilon}_{1}))$.
Since we treat polygons as equivalent under scaling, translations, rotations, reflections, or any combination of these transformations, we define two vectors $\boldsymbol{\varepsilon}_{1},\boldsymbol{\varepsilon}_{2}\in\{\pm 1\}^{m}$ to be equivalent if one can be obtained from the other by applying a finite sequence of the shifting operator $\mathcal{S}$ and the reflection operator $\mathcal{R}$. A simple calculation shows that ${\mathcal{R}}^{2}={\mathcal{S}}^{2m}={\rm Id}$, where $\rm Id$} denotes the identity operator. Furthermore, we have ${\mathcal{S}}^{k}{\mathcal{R}}={\mathcal{R}}{\mathcal{S}}^{2m-k}$ for any integer $k\in[1,2m]$. Hence, $\boldsymbol{\varepsilon}_{1}$ is equivalent to $\boldsymbol{\varepsilon}_{2}$ if and only if ${\mathcal{R}}^{k_{1}}{\mathcal{S}}^{k_{2}}(\boldsymbol{\varepsilon}_{1})=\boldsymbol{\varepsilon}_{2}$ for some $k_{1}\in\{0,1\}$ and $k_{2}\in\{0,1,\ldots,2m-1\}$. Under this convention, the number of distinct solutions to Problem 2.1 equals that of Problem 4.1.
Remark 4.2.
Let $m\geq 3$ be an integer with an odd factor. Recall the brute-force algorithm mentioned in Remark 1.2. In fact, by Theorem 4.1, it can also be used to generate all optimal polygons in ${\mathcal{E}}_{m}$, up to scaling, translation, rotation, and reflection. We again take $m=12$ as an example. After checking all vectors $\boldsymbol{\varepsilon}\in\{\pm 1\}^{12}$ for which (4.1) holds, by Theorem 4.1, we identify all optimal polygons in ${\mathcal{E}}_{12}$ (up to scaling, translation, rotation and reflection), shown in Figure 2 (a1) and (a2). Note that this algorithm for constructing all optimal polygons in ${\mathcal{E}}_{m}$ differs from the one proposed in [25]. A detailed comparison of the efficiency of the two algorithms is left for future work.
@@ 4.1. Characterization of optimal polygons
With the help of Theorem 4.1, we give a proof of Theorem 1.2.
@@
Combining with Lemma 2.1, we derive that
4.3. Proof of Theorem 4.1
In this subsection we prove Theorem 4.1. We begin by recalling the definition of Reuleaux polygons, which will be used in the sequel.
@@
Let $P$ be an optimal convex $m$-gon, where $m\geq 3$ is an integer with an odd factor. By Theorem 4.3 and Theorem 4.2 (i), $P$ is inscribed in a Reuleaux polygon $R$ with $r$ boundary arcs, where $r$ is an odd integer with $3\leq r\leq m$. Label the vertices of $R$ as $\boldsymbol{w}_{1},\dotsc,\boldsymbol{w}_{r}$ in counterclockwise order, and extend the indexing cyclically by setting $\boldsymbol{w}_{r+i}=\boldsymbol{w}_{i}$ for each $i\in[r]$. Define $h=\frac{r-1}{2}$. Then for each $i\in[r]$, the vertex $\boldsymbol{w}_{i}$ is the the center of the arc ¿ $\boldsymbol{w}_{{i+h}}\boldsymbol{w}_{{i+h+1}}$ on the boundary of $R$. This can be seen as follows. By Theorem 4.2 (ii), we can start at $\boldsymbol{s}_{1}:=\boldsymbol{w}_{1}$ and construct a star polygon $S$ by sequentially connecting pairs of vertices in $R$ at maximal distance in the counterclockwise direction. Let $\boldsymbol{s}_{i}$ denote the $i$-th vertex of $R$ visited during this construction. According to [26, Page 256], the vertices $\boldsymbol{s}_{1},\boldsymbol{s}_{3},\dotsc,\boldsymbol{s}_{r},\boldsymbol{s}_{2},\boldsymbol{s}_{4},\dotsc,\boldsymbol{s}_{r-1}$ are ordered counterclockwise. Comparing this with the labeling $\boldsymbol{w}_{1},\ldots,\boldsymbol{w}_{r}$, we have
@@
Figure 3. One of the optimal dodecagons as in Figure 2 (a2), with vertices $\boldsymbol{v}_{1},\dotsc,\boldsymbol{v}_{12}$. The corresponding Reuleaux polygon consists of $9$ circular arcs with vertices $\boldsymbol{v}_{c_{1}},\dotsc,\boldsymbol{v}_{c_{9}}$ indicated in red dots. The dashed lines connect pairs of vertices at maximal distance.
@@
Assume that $P\in{\mathcal{E}}_{m}$. It follows from Theorem 4.3 that $P$ is equilateral. We now prove that $P$ satisfies (4.5). Assume without loss of generality that $\boldsymbol{e}_{1},\dotsc,\boldsymbol{e}_{m}$ are arranged counterclockwise. Let $\boldsymbol{v}_{1},\dotsc,\boldsymbol{v}_{m}$ be the vertices of $P$ such that $\boldsymbol{e}_{j}=\boldsymbol{v}_{j+1}-\boldsymbol{v}_{j}$ for each $j\in[m-1]$ and $\boldsymbol{e}_{m}=\boldsymbol{v}_{1}-\boldsymbol{v}_{m}$. For each $i\in[m]$, we denote $\boldsymbol{v}_{m+i}=\boldsymbol{v}_{i}$, and let $\theta_{i}$ be the counterclockwise angle from $\boldsymbol{e}_{i-1}$ to $\boldsymbol{e}_{i}$, where $\boldsymbol{e}_{0}:=\boldsymbol{e}_{m}$. Then we have $\alpha_{1}=0$ and $\alpha_{i}=\alpha_{i-1}+\theta_{i}$ for each $i\in\{2,\ldots,m\}$. Moreover, since $P$ is strictly convex, we have $\theta_{i}\in(0,\pi)$, for $i\in[m]$.